OpenStudy (anonymous):

Rewrite the standard form equation below into slope-intercept form. x – 2y = 3

OpenStudy (anonymous):

Just make 'y' the subject!

OpenStudy (anonymous):

Slope = 1.000/2.000 = 0.500 x-intercept = 3/1 = 3.00000 y-intercept = 3/-2 = -1.50000